Compare the following rates. Label as fastest, medium, and slowest: 1780 yards per minute, 62 miles per hour, or 86 feet per second?
Found 2 solutions by mananth, avrybrdly93:
Answer by mananth(16946)


Compare the following rates. Label as fastest, medium, and slowest:


1780 yards per minute,
= 1780*3/60
=89ft/s



62 miles per hour,
=62*5280/3600
=89.9 ft/s


or 86 feet per second?
86ft/s
You decide

Answer by avrybrdly93(1)

For each of these numbers, we need to convert them so that they have the same units. For example, in order to correctly compare 1780 yards per minute with 62 miles per hour, both numbers should be in either miles per hour or yards per minute. This can easily be done by just a couple conversions. To make things easier to understand, lets call 1780 yards per minute speed A, 62 miles per hour speed B, and 86 feet per second speed C.
First of all, we know that there are 1760 yards in one mile and 60 minutes in one hour. All we need to do is divide 1780 yards into 1760 yards. This will tell us how many miles per minute 1780 yards per minute is. After this, all we need to do is convert minutes to hours. Because of this, 1.0113 (this is speed A = 1780/1760 in miles per minute) becomes 68.6818 miles per hour. We are able to see that speed A is greater than speed B (68.6818 > 62). Now, all that needs to be done is converting 86 feet per second to miles per hour and then compare it to speed A and B.
*Hint: 5280 feet = 1 mile
3600 secs = 1 hour